# Lean Workbook: A large-scale Lean problem set formalized from natural language math problems

### v2

@@ 1 Introduction
In this work, we present Lean Workbook: an iterative autoformalization pipeline, together with a large-scale Lean problem set. We train our autoformalization model based on active learning. At each turn, we use our model to translate natural language problems into formal statements and back-translate to natural language problems collected from the math contest forum [^1] https://artofproblemsolving.com/community. We use Lean compiler and Natural Language Inference (NLI) to check if it is a valid formalization. We sample invalid formalization and require human experts to modify them into a valid formalization and add them to the training set. Through the supplement of human-labeled data pairs, the translation model gradually learned to translate between Lean 4 formal language and natural language questions of different types of problems.
We autoformalized 57K math problems in the final round. Manual examination reports an accuracy of 93.5% of a random sample of the Lean Workbook. The same filtering process produces 21 new formal statements of the IMO questions which do not appear in Compfiles  [^2] https://github.com/dwrensha/compfiles.
@@ 2 Preliminaries
Formal proof involves establishing claims that are expressed in precise mathematical terms in programming languages. Lean 4, which is the latest version of Lean Theorem Prover, aims to provide an open-source platform for correct and maintainable code for formal verification. The Lean language can claim a theorem and prove it by tactics or pretend to complete the proof using "sorry". Lean 4 will return a "No goals" signal if the proof is completed.
@@ 4.1 First-round pipeline
The first-round data collection is fed to our translate model, which is initialized from InternLM-Math-Plus-20B 2023internlm which has been pre-trained on Lean-related datasets. The model is fine-tuned for three epochs with a learning rate of $54e-$ 5, with two different but fixed prompts for each translation direction. This translated data serves as a starting point for further iteration. Fine-tuning uses 32 A100 GPUs and can be finished within several hours.
@@
The well-defined subset contains 6652 different tags in total, with 223 tags containing over 100 samples. These tags cover a large range of questions from contest-level knowledge points to high-school courses. More than three-thirds of the samples are labeled with algebra-relevant tags, while geometry-related tags are rarely witnessed. It is also noticed that some tags are wrong, especially the "number theory" tag is often allocated to inequality problems. Following these findings, we will remain keeping working with tags over 100 samples in later analysis and will pay special attention to wrong tags during manual diagnostic.
@@ 4.2 Data Diagnostic and Iteration pipeline
Compiling Correctness test To ensure the accuracy of the formal statements produced by our translation pipeline, each translated formal theorem undergoes a correctness check within a Lean 4 environment. Initially, the theorem statements are verified independently, using a placeholder "by sorry" for the proofs, to filter out incorrect statements in advance. The complete theorem, including proofs, is then examined. The major bottleneck of this step is the compiling cost of Lean 4 projects. To facilitate the process, we build up a Lean 4 read-eval-print loop (REPL), utilizing Lean 4’s runtime meta-programming facility, which allows for the verification of Lean 4 statements in an interpreted mode. The correctness test program can be executed in a multi-process style and can be finished in one hour with a 32-core CPU. Our test environment is based on Lean v4.7.0 with Mathlib4 of the same version (which can be cloned by specifying the tag v4.7.0). The Minif2f environment is equivalent to the header in https://huggingface.co/datasets/internlm/Lean-Workbook/discussions/1.
@@
Diagnostic and Human labeling Diagnostic for the data mainly focuses on two kinds of samples: the ones that do not pass the compiling correctness test and the ones that pass the test but do not prove to be a correct translation with a positive NLI feedback. The other samples that pass the NLI test are considered to be correct for now. In the first three rounds of our iterations, these two kinds of samples both have relatively obvious patterns. Thus we conclude and modify them accordingly with three human experts who are familiar with both Lean and contest-level math problems  [^6] They all won a prize in the National Mathematical Olympiad Contest..
@@ 5.3 Effectiveness and discussion
For an intuitive comparison of the effectiveness of our active learning pipeline, we derive the CPN and NPN for three models: The first-round model which is used for the initial filtering, the final-round model generating our dataset, and the final model further fine-tuned on our dataset Lean Workbook. The results are shown in Table 3.
@@
This table clearly shows the effectiveness both of our pipeline and our dataset. The human-labeled data and filtered dataset achieve a gain of over 20000 more correct samples for both the compile test and the NLI test, which promisingly indicates that this form of active learning can be further iteratively utilized.

### v3

@@ 1 Introduction
I do believe that problems are the heart of mathematics.P. R. Halmos
@@
In this work, we present Lean Workbook: an iterative autoformalization pipeline, together with a large-scale Lean problem set. We train our autoformalization model based on active learning. At each turn, we use our model to translate natural language problems into formal statements and back-translate to natural language problems collected from the math contest forum [^1] https://artofproblemsolving.com/community. We use Lean compiler and Natural Language Inference (NLI) to check if it is a valid formalization. We sample invalid formalization and require human experts to modify them into a valid formalization and add them to the training set. Through the supplement of human-labeled data pairs, the translation model gradually learned to translate between Lean 4 formal language and natural language questions of different types of problems. We autoformalized 57K math problems in the final round. Manual examination reports an accuracy of 93.5% of a random sample of the Lean Workbook. The same filtering process produces 21 new formal statements of the IMO questions which do not appear in Compfiles  [^2] https://github.com/dwrensha/compfiles.
@@ 2 Preliminaries
Formal proof involves establishing claims that are expressed in precise mathematical terms in programming languages. Lean 4, which is the latest version of Lean Theorem Prover, aims to provide an open-source platform for correct and maintainable code for formal verification. In the Lean language, users can define a theorem and prove it by tactics or pretend to complete the proof using "sorry". Lean 4 will return a "No goals" signal if the proof is completed.
@@ 3.2 Automatic Theorem Proving
3.3 Data curation
In the field of formal language, several works have established their methods of development and curation of datasets. Runtime information has proved to be beneficial for the construction of datasets containing whole proofs, which can be extracted via tools such as LeanDojo and CoqGym(yang2023leandojo, ; yang2019coqgym). Meanwhile, human annotations and human interaction with language models can provide valuable assistance towards stepwise proofs(song2024towards; welleck2023llmstep, ). Aside from large-scale datasets, there are also highly accurate benchmarks for formally verifying proofs(lohn2024minicodeprops; zheng2021minif2f, ). Our work shares a similarity with these works in that human experts and LLMs contribute together to the autoformalization data construction.
@@ 4.1 First-round pipeline
The first-round data collection is fed to our translation model, which is initialized from InternLM-Math-Plus-20B 2023internlm which has been pre-trained on Lean-related datasets. The model is fine-tuned for three epochs with a learning rate of $54e-$ 5, with two different but fixed prompts for each translation direction. This translated data serves as a starting point for further iteration. Fine-tuning uses 32 A100 GPUs and can be finished within several hours.
@@
The well-defined subset contains 6652 different tags in total, with 223 tags containing over 100 samples. These tags cover a large range of questions from contest-level knowledge points to high-school courses. More than three-fourths of the samples are labeled with algebra-relevant tags, while geometry-related tags are rarely witnessed. It is also noticed that some tags are wrong, especially the "number theory" tag is often allocated to inequality problems. Following these findings, we will remain keeping working with tags over 100 samples in later analysis and will pay special attention to wrong tags during manual diagnostic.
@@ 4.2 Data Diagnostic and Iteration pipeline
Compiling Correctness test To ensure the accuracy of the formal statements produced by our translation pipeline, each translated formal theorem undergoes a correctness check within a Lean 4 environment. Initially, the theorem statements are verified independently, using a placeholder "by sorry" for the proofs, to filter out incorrect statements in advance. The complete theorem, including proofs, is then examined. The major bottleneck of this step is the compiling cost of Lean 4 projects. To facilitate the process, we build up a Lean 4 read-eval-print loop (REPL), utilizing Lean 4’s runtime meta-programming facility, which allows for the verification of Lean 4 statements in an interpreted mode. The correctness test program can be executed in a multi-process style and can be finished in one hour with a 32-core CPU. Our test environment is based on Lean v4.8.0-rc1 with Mathlib4 of the same version (which can be cloned by specifying the tag v4.8.0-rc1).
@@
Diagnostic and Human labeling Diagnostic for the data mainly focuses on two kinds of samples: the ones that do not pass the compiling correctness test and the ones that pass the test but do not prove to be a correct translation with a positive NLI feedback. The other samples that pass the NLI test are considered to be correct for now. In the first three rounds of our iterations, these two kinds of samples both have relatively obvious patterns. Thus we conclude and modify them accordingly with three human experts who are familiar with both Lean and contest-level math problems  [^6] They all won a prize in the National Mathematical Olympiad Contest.. Each evaluator was assigned an equal number of problems, ensuring a balanced distribution of the workload. On average, each problem required approximately two to five minutes for evaluation.
@@ 5.3 Effectiveness and discussion
For an intuitive comparison of the effectiveness of our active learning pipeline, we derive the CPN and NPN for three models: The first-round model which is used for the initial filtering, the final-round model generating our dataset, and the final model further fine-tuned on our dataset Lean Workbook. The results are shown in Table 3. We also listed the accuracy of MiniF2F valid and test set when an InternLM2-Math-Plus model is fine-tuned on MiniF2F with and without our Lean-Workbook dataset in Table 4.
@@
Table 4: We report the accuracy for an InternLM2-Math-Plus model fine-tuned on MiniF2F only and with our Lean-Workbook dataset.
Train Dataset MiniF2F-valid Acc. MiniF2F-test Acc.
Mathlib 44,3 37.3
+ Lean Workbook 50.4 46.7
This table clearly shows the effectiveness both of our pipeline and our dataset. The human-labeled data and filtered dataset achieve a gain of over 20000 more correct samples for both the compile test and the NLI test, which promisingly indicates that this form of active learning can be further iteratively utilized. The increase in MiniF2F accuracy also demonstrates a significant improvement in performance when using our extended dataset.
@@ 6 Conclusion
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory, and funded by the project JF-P23KK00072-2-DF.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See abstract and introduction
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section Limitations
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Proving math theorems do not have apparent negative social impacts.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We specify how we train translation models in Section 4.1.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A] We only fine-tune once for each iteration.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Details are described in Section 4.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We cite MiniF2F and Proofnet.
    2. (b)
      Did you mention the license of the assets? [Yes] We report in Section 4.1.
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] Math problems do not contain PII.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]